OpenStudy (philcua):

What value of y makes the equation true? (–9.7 • 24) • (–0.25) = –9.7 • (24 • y) y = ?

OpenStudy (philcua):

Can you help me with the last question for test @midhun.madhu1987

OpenStudy (philcua):

Can you help @midhun.madhu1987

OpenStudy (philcua):

@bibby

OpenStudy (bibby):

is that multiplication

OpenStudy (philcua):

Yes

OpenStudy (bibby):

do the multiplications then –9.7 * 24 * –0.25

OpenStudy (philcua):

I am supposed to figure out what y is openstudy kepps putting question marks

OpenStudy (bibby):

if you refresh it should work fine

OpenStudy (philcua):

Yes that equals -232

OpenStudy (philcua):

I think

OpenStudy (bibby):

58.2 use a calculator or learn to multiply decimals

OpenStudy (philcua):

Yes but how would you figure out what y means

OpenStudy (bibby):

algebra...

OpenStudy (bibby):

that's not even necessary actually

OpenStudy (philcua):

Yes but I kneed the value of y

OpenStudy (bibby):

THAT'S WHAT ALGEBRA IS FOR

OpenStudy (bibby):

(–9.7 * 24) * (–0.25) = –9.7 * (24 * y) (–9.7 * 24) * (–0.25) = (–9.7 * 24) * y does that help?

OpenStudy (philcua):

Yes but how am I supposed to like find the value of y I have tried different numbers already

OpenStudy (bibby):

\(\cancel{(–9.7 * 24)} * (–0.25) =\cancel{(–9.7 * 24)}* y\)

OpenStudy (bibby):

flutter openstudy

OpenStudy (bibby):

\(\cancel{(–9.7 * 24)}*-0.25=\cancel{(–9.7 * 24)}*y=\)

OpenStudy (bibby):

refresh

OpenStudy (philcua):

so y would equal -0.25

OpenStudy (bibby):

yeah

OpenStudy (philcua):

Ok thank you so much but are you sure that is the answer

OpenStudy (bibby):

yeah

OpenStudy (bibby):

not necessary

OpenStudy (philcua):

got it right thank you bro